When striking, the pike, a predatory fish, can accelerate from rest to a speed of 3.6 m/s in 0.13 s.

How far does the pike move during this strike?
Answer #1

a= v/t = 3.6/0.13 = 27.7 m/s2

S= 0.5at2

S= 0.5x27.7x0.132

S= 0.234 m

-------------------------------